# Self-Routing RAG: Binding Selective Retrieval with Knowledge Verbalization

### v1

@@ C.1 Self-Routing RAG: Hyperparameters
In Table 6, we present the F1 score and AUROC using Llama-2-7B-Chat as the LLM. Correctness is defined as preferring self-verbalized knowledge when it is better than or equal to performing retrieval in terms of the downstream question answering performance. Compared to three baseline methods, the heuristics SR-RAG uses achieves the best F1 and AUROC over the short-form and closed-set subset of the training set. Building upon this paper’s results, future work can further study leveraging more advanced uncertainty quantification methods (Wu et al., 2024b; Moskvoretskii et al., 2025a) to build more accurate source selection labels.
@@
To study the influence of the kNN policy datastore size on the performance of SR-RAG, we randomly sample a subset of the training set (50k samples in total) to construct the datastore. As shown in Figure 6 (blue), reducing the kNN policy datastore to 25k (half) has minimal impact. Even with just 1k examples, performance drops only slightly, suggesting potential for lightweight and memory-efficient deployments.
@@
The number of neighbors retrieved from the datastore determines the localness of the focus in the hidden representation space. While task-specific neighborhood size can be tuned, SR-RAG generally works well with 10 to 30 neighboring data points (Figure 6 middle), indicating its robustness. However, using a too small (e.g., 5) or a too large (e.g., more than 50) neighborhood harms the performance of SR-RAG.
@@
We further study the impact of layer selection on SR-RAG’s performance in Figure 6 part 3 (green). Overall, we observe that using a middle layer’s hidden states has more superior performance over using the last layer’s. In the next section, we visualize the hidden state space across LLMs and show that middle layers tend to learn better representations for the model uncertainty as well as task types.
C.2 Hidden State Space of the Source Reflection Token
@@
In Figure 7, we visualize <EOQ> hidden states on PopQA using t-SNE. Clear separation emerges between instances where self-verbalized knowledge is helpful (green) and unhelpful (red). Middle layers show stronger clustering than final layers, supporting the use of these representations for kNN-based source selection.
@@
We further randomly sample 500 data points from the SR-RAG train set and visualize their hidden states. Figure 8 shows that SR-RAG also learns to cluster instances by task type (e.g., fact-checking, closed QA, long-form generation). This suggests that source reflection tokens carry semantically meaningful and task-sensitive representations on top of uncertainty information. These behaviors serve as the foundation of our nearest neighbor-based source selection approach.
C.3 Can SR-RAG learn more from self-generated knowledge?
In SR-RAG training, self-verbalized knowledge contexts from the LLM are utilized to fine-tune the model via behavior cloning (stage 1) or preference alignment (stage 2). Alternatively, is it possible that the model can learn to verbalize even higher-quality knowledge from off-policy samples generated by a stronger LLM? We investigate this hypothesis by running GenRead-based knowledge verbalization on Llama-3.3-70B-Instruct and directly use the knowledge to prepare SR-RAG training data with Llama-2-7B-Chat. Table 7 compares the performance of fine-tuning Llama-2-7B-Chat with this data versus using self-generated knowledge for SR-RAG. Interestingly, using self-generated positive and negative knowledge pairs is more beneficial for unlocking the model’s ability to generate useful knowledge, outperforming the alternative by 4.9% higher answer accuracy and 5.0% higher win rate over retrieval. It is also possible that the ability of the teacher model needs to be close to the fine-tuned model for its knowledge to be useful. We leave this investigation to future work.
@@ C.3 Can SR-RAG learn more from self-generated knowledge?
Verbalization Performance ()($\uparrow$)( ↑ )
@@
C.4 Qualitative Study
@@
- What was Mr. Smith’s first name in Mr. Smith goes to Washington? (<Self> $>$ <Wiki>)
- Who was the original singer of Knocking on Heaven’s Door? (<Self> $>$ <Wiki>)
- When did sir Arthur Conan Doyle wrote his first book? (<Wiki> $>$ <Self>)
@@
- The battle of Hastings in 1066 was fought in which country? (<Self> $>$ <Wiki>)
- When Belgium declared its independence in 1830 it broke away from ___ control? (<Self> $>$ <Wiki>)
- When did the French come to the new world? (<Self> $>$ <Wiki>)
@@
In Figure 9 and Figure 10, we show two qualitative examples from TriviaQA with the fine-tuned Llama-2-7B-Chat as the backbone LLM. For both examples, we show the LLM’s prediction conditioned on both retrieved knowledge and the verbalized knowledge. While the verbalized knowledge is more concise than retrieved paragraphs, it better identifies the correct answers for both questions. This is because the core information (i.e., ”first name”, ”island”) only takes up a small portion of the query. As a result, dense retrieval is noise-prone, bringing in only distantly relevant information. By contrast, the LLM as a knowledge generator dynamically adjusts its focus, guiding the knowledge generation to include the core information. As a result, the LLM can correctly answer both questions with self-verbalized knowledge but provides incorrect answers with the low-quality retrieved knowledge. In the source selection process, the LLM directly predicts the source preference $p_{M}(\text{{\textless s\textgreater}}|q)$. In addition, questions from the train set sharing the same topic and difficulty level are retrieved to form the source preference distribution $p_{D}(\text{{\textless s\textgreater}}|q)$. In the first example (Figure 9), both distributions prefer the LLM itself as the knowledge source. By contrast, in the second example (Figure 10), $p_{M}(\text{{\textless s\textgreater}}|q)$ has a high uncertainty and $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source distribution to select the LLM itself as the knowledge source.

### v2

@@ D.1 Self-Routing RAG: Hyperparameters
In Table˜7, we present the F1 score and AUROC using Llama-2-7B-Chat as the LLM. Correctness is defined as preferring self-verbalized knowledge when it is better than or equal to performing retrieval in terms of the downstream question answering performance. Compared to three baseline methods, the heuristics SR-RAG uses achieves the best F1 and AUROC over the short-form and closed-set subset of the training set. Building upon this paper’s results, future work can further study leveraging more advanced uncertainty quantification methods (Wu et al., 2024b; Moskvoretskii et al., 2025b) to build more accurate source selection labels.
@@
To study the influence of the kNN policy datastore size on the performance of SR-RAG, we randomly sample a subset of the training set (50k samples in total) to construct the datastore. As shown in Figure˜6 (blue), reducing the kNN policy datastore to 25k (half) has minimal impact. Even with just 1k examples, performance drops only slightly, suggesting potential for lightweight and memory-efficient deployments.
@@
The number of neighbors retrieved from the datastore determines the localness of the focus in the hidden representation space. While task-specific neighborhood size can be tuned, SR-RAG generally works well with 10 to 30 neighboring data points (Figure˜6 middle), indicating its robustness. However, using a too small (e.g., 5) or a too large (e.g., more than 50) neighborhood harms the performance of SR-RAG.
@@
We further study the impact of layer selection on SR-RAG’s performance in Figure˜6 part 3 (green). Overall, we observe that using a middle layer’s hidden states has more superior performance over using the last layer’s. In the next section, we visualize the hidden state space across LLMs and show that middle layers tend to learn better representations for the model uncertainty as well as task types.
D.2 Hidden State Space of the Source Reflection Token
@@
In Figure˜9, we visualize <EOQ> hidden states on PopQA using t-SNE. Clear separation emerges between instances where self-verbalized knowledge is helpful (green) and unhelpful (red). Middle layers show stronger clustering than final layers, supporting the use of these representations for kNN-based source selection.
@@
We further randomly sample 500 data points from the SR-RAG train set and visualize their hidden states. Figure˜10 shows that SR-RAG also learns to cluster instances by task type (e.g., fact-checking, closed QA, long-form generation). This suggests that source reflection tokens carry semantically meaningful and task-sensitive representations on top of uncertainty information. These behaviors serve as the foundation of our nearest neighbor-based source selection approach.
@@
Verbalization Performance $(\uparrow)$
@@
D.3 Can SR-RAG learn more from self-generated knowledge?
In SR-RAG training, self-verbalized knowledge contexts from the LLM are utilized to fine-tune the model via behavior cloning (stage 1) or preference alignment (stage 2). Alternatively, is it possible that the model can learn to verbalize even higher-quality knowledge from off-policy samples generated by a stronger LLM? We investigate this hypothesis by running GenRead-based knowledge verbalization on Llama-3.3-70B-Instruct and directly use the knowledge to prepare SR-RAG training data with Llama-2-7B-Chat. Table˜8 compares the performance of fine-tuning Llama-2-7B-Chat with this data versus using self-generated knowledge for SR-RAG. Interestingly, using self-generated positive and negative knowledge pairs is more beneficial for unlocking the model’s ability to generate useful knowledge, outperforming the alternative by 4.9% higher answer accuracy and 5.0% higher win rate over retrieval. It is also possible that the ability of the teacher model needs to be close to the fine-tuned model for its knowledge to be useful. We leave this investigation to future work.
D.4 Qualitative Study
In Figure˜7 and Figure˜8, we show two qualitative examples from TriviaQA with the fine-tuned Llama-2-7B-Chat as the backbone LLM. For both examples, we show the LLM’s prediction conditioned on both retrieved knowledge and the verbalized knowledge. While the verbalized knowledge is more concise than retrieved paragraphs, it better identifies the correct answers for both questions. This is because the core information (i.e., "first name", "island") only takes up a small portion of the query. As a result, dense retrieval is noise-prone, bringing in only distantly relevant information. By contrast, the LLM as a knowledge generator dynamically adjusts its focus, guiding the knowledge generation to include the core information. As a result, the LLM can correctly answer both questions with self-verbalized knowledge but provides incorrect answers with the low-quality retrieved knowledge. In the source selection process, the LLM directly predicts the source preference $p_{M}(\text{{\textless s\textgreater}}|q)$. In addition, questions from the train set sharing the same topic and difficulty level are retrieved to form the source preference distribution $p_{D}(\text{{\textless s\textgreater}}|q)$. In the first example (Figure˜7), both distributions prefer the LLM itself as the knowledge source. By contrast, in the second example (Figure˜8), $p_{M}(\text{{\textless s\textgreater}}|q)$ has a high uncertainty and $p_{D}(\text{{\textless s\textgreater}}|q)$ corrects the source distribution to select the LLM itself as the knowledge source.
@@ D.4 Qualitative Study
- What was Mr. Smith’s first name in Mr. Smith goes to Washington? (<Self> >> <Wiki>)
- Who was the original singer of Knocking on Heaven’s Door? (<Self> >> <Wiki>)
- When did sir Arthur Conan Doyle wrote his first book? (<Wiki> >> <Self>)
@@
- The battle of Hastings in 1066 was fought in which country? (<Self> >> <Wiki>)
- When Belgium declared its independence in 1830 it broke away from ___ control? (<Self> >> <Wiki>)
- When did the French come to the new world? (<Self> >> <Wiki>)